Question 5 (20 points): Calculate the ENPV for buying and drilling an oil lease assuming a minimum ROR of 12.5% and the following estimated costs and success/falure rates The lease costs $400,000 dollars at time zero and drilling will start at year 1 with a cost of $150,000 dollars There is a 100% lease probability of drilling a well ifwe decade to purchase the Success Cases. If drilling is successful, then there are two possibulities 1. A 50% probability that the well yields an annual income of $250,000 dollars for 12 years (from year 2 to year 13) with an environmental remediation cost of $100,000 in year 13 A 50% probability that the well yields annual income of$200,000 dollars for 9 years 2 (from year 2 to year 10) with an environmental remediation cost of $S0,000 im year 10 Is this investment economically satisfactory? Explam your work in detail including all the required factors and calculations XLL
Answer #1

Present value of the costs of leasing and Drilling = 400,000 + 150,000/(1+12.5%) = -533,333.33

With failure PV = -75000/(1+12.5%) = 66,666.66

With actual 20% (40% *50%=20%)probability of succes in case 1 for year 0 = Annuity * (1-(1+r)-n)/r *1/(1+r) - envornmental cost /(1+r)13 = 250000 * (1-(1+12.5%)-12)/12.5% *1/(1+12.5%) - 100,000/(1+12.5%)13​ = 1323588.8917


With actual 20% (40% *50%=20%)probability of succes in case 2 for year 0 = Annuty * (1-(1+r)-n)/r *1/(1+r) + envornmental cost /(1+r)13 = 250000 * (1-(1+12.5%)-12)/12.5% *1/(1+12.5%) + 100,000/(1+12.5%)13​ =914111.0786

ENPV = - 533,333.33 + 0.6 *(66,666.66) + 0.2* 1323588.8917 + 0.2 *  914111.0786 = -125793.34

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well



Since the ENPV of the roject is negative the project should not be started.